OpenStudy (anonymous):

what is an equation for the translation of (x-2)^2 + (y+1)^2=16 by 3 units left and 5 units up?

OpenStudy (anonymous):

is it (x+1)^2 +(y-4)^2=16

OpenStudy (anonymous):

3 left-- x=0 becomes x'= -3=> x'=x-3=> x=x'+3 5 up--- y=0becomes y'=5 => y'=y+5=> y=y'-5 cubstitte for x & y

OpenStudy (anonymous):

why are u setting x and y = to 0? are you just ignoring the equation (x-2)^2 and making it (x+3)^2

OpenStudy (anonymous):

+ (y-5)^2

OpenStudy (anonymous):

i did not set anything.. for translation, I was merely showing the translation of one of the points.... set x=x'+3 and y=y'-5 in original, the new co-ordinate system of x' and y' is obtained wherein, X'Y' plane is the translated XY plane

OpenStudy (anonymous):

this metjod is used in what we call as "computer vision" where points need to translated, rotated, skewed, etc in 3D and find how the 3D world transforms

OpenStudy (anonymous):

so, what you got is correct, just more elaborate as t

OpenStudy (anonymous):

wh

OpenStudy (anonymous):

o why

OpenStudy (anonymous):

thanks